A 5.0-kg object is pulled along a horizontal surfa
Answer #1

1. W = F.d = |F||d| cos@

@ is the angle between F and d.

W = 11.5 * 7.4 * cos30 = 73.7 J
ANs(a)


2. W = 12 * 4.15 * cos15 = 48.1 J

ans(D)

4. In vertical,

N + Fsin30 - mg = 0

N = 137.5 - (15 sin30) = 130 N


in horizontal,

f - Fcos30 = 0

f = friction force = uk N

u N = F cos30

130 uk = 15 cos30

uk = 0.10


4. in vertical,

Fnet= N + 16sin37 - mg = 0

N = (2 x 9.8) - (16 sin37) = 9.97 N


friction force, f =uk N = 0.50 x 9.97 = 5 N

work done = f.dcos180 = 5 * 7 *cos180 = - 35 J